OpenStudy (anonymous):

what is the slope intercept form of 4x=7

OpenStudy (hanner_b_nanner):

y=4x+7...

OpenStudy (anonymous):

ok thank u.

OpenStudy (anonymous):

how did u get that

OpenStudy (hanner_b_nanner):

slop formula is y=mx+b, 4 is the m and 7 is the b

OpenStudy (anonymous):

ok can u do this one more for me x-3y=8

OpenStudy (anonymous):

if the equation is 9x+5y=-3 is the answer 5y-9x+~3

OpenStudy (hanner_b_nanner):

First one, y=1/3x-8/3. That's what http://hotmath.com said....

OpenStudy (anonymous):

ok is the second on right